What is the difference between the sum of all postive odd integers less than 102 and the sum of all positive even integers less than 102 ?
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

What is the difference between the sum of all postive odd integers less than 102 and the sum of all positive even integers less than 102 ?
==========
--> 1 to 101
-------------
The sum of 51 odd numbers = 51^2 = 2601
-------
The sum of all numbers from 1 to 101 = 5151
--------
Evens = 5151 - 2601 = 2550

Answer by Edwin McCravy(20055)

 
The other tutor's answer is incorrect.  

The first 101 integers consist of 50 even numbers and 51 odd numbers

sum of odds  = 1 + 3 + 5 + 7 + ··· + 95 + 97 + 99 + 101
sum of evens =     2 + 4 + 6 + ··· + 94 + 96 + 98 + 100
-------------------------------------------------------
difference   = 1 + 1 + 1 + 1 + ··· +  1 +  1 +  1 +   1

When we subtracted them above, term by term, there is a 1 
in each of the positions of the 51 odd numbers. The sum of 
51 1's is 51.

Answer: 51

Edwin